eljaja:

can someone help me understand and find the domain and range of f(x) = sec 2x. Im not really familiar with domain and range at all so you can explain like what they are and how im supposed to find them

lowkey:

Are you allowed to use a graphing calculator?

Shadow:

^

lowkey:

Cause that answer depends on what I'm about to teach you lol

eljaja:

Yes lol

sillybilly123:

looking at a graph first is not a bad way to do it

eljaja:

Okay... I put it into desmos. What do I exactly do now?

sillybilly123:

maybe look at the graph? for all inputs [values of "x"], you are getting something defined out of that "function",... , apart from the bits when sec(x) = 1/cos(x) = 1/0 so your domain has to exclude the singularities?

sillybilly123:

do you know what a function is defined as? you can glue these ideas of input/output ---- domain/range by know what that means. https://www.mathsisfun.com/sets/function.html

eljaja:

Yes I obviously looked at the graph. I think I see what you mean. Is the domain going to infinity or no?

lowkey:

Gradually but there are domain restrictions because the graph is not continuous.

sillybilly123:

\(\checkmark\)

eljaja:

Like exactly what kind of restrictions? how can I tell that?

sillybilly123:

Put the horse before the cart If \(sec(x)\) goes off into \(\infty\) for certain values of \(x\), then those values that do so are excluded from the definition of \(\sec(x)\) as a function

sillybilly123:

so it's Domain is limited

eljaja:

um

sillybilly123:

to which number does \(\sec(\pi/2)\) map to?

sillybilly123:

so is \(\pi/2\) within the domain of the function \(f(x) = \sec(x)\)?

eljaja:

Isn't it undefined and doesn't map to any number or no?

sillybilly123:

yes, if a "plausible" function maps \(f(x_o) \to \infty\) then there is a problem with this "function" at \(x = x_o\), a singularity. So we must exclude the various \(x_o\)'s from the Domain Once we do, the range bit is easy

eljaja:

Ok i see and how do we exclude?

sillybilly123:

this is stuff i copied for sec(x), so adjust for sec(2x) <-- doubled-frequency  

eljaja:

Im sorry I have to go. Thanks for helping, or at least trying to help. I might have to find someone else later that can like explain it a lot better. Thanks anyways.

sillybilly123:

Trust me, when the penny drops, this is the explanation that you will remember :) Happy hunting !!

eljaja:

Sorry its not.... i just started learning domain and range because im in 9th grade. Your explanation is just too complex for someone that just started learning this.